OpenStudy (anonymous):

What is the volume of a solid obtained by rotating the region between the graph of y = (1/2)sin^2 (x^2) and the x-axis for 0 ≤x≤sqrtpi about the y-axis?

OpenStudy (anonymous):

The main problem I have is that I can't solve for x.

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (callisto):

Put y =0 0 = (1/2)sin^2 (x^2) sin^2 (x^2) = 0 sin (x^2) = 0 x^2 =...???

OpenStudy (anonymous):

0?

OpenStudy (callisto):

and??

OpenStudy (anonymous):

0 is all I can think of.

OpenStudy (callisto):